A factory can make ice cream or popsicles. In the course of 1 day, the factory can either make 1,000 gallons of ice cream or 3,000 popsicles. The opportunity cost of 1 gallon of ice cream is:

Question 11 options:

1/3 of a popsicle.

Zero popsicles--just make both at the same time.

3 popsicles.

1 popsicle.

Decision makers in companies should concern themselves with:

Question 12 options:

maximizing the company's value over the entire life of the company.

maximizing revenue

maximizing profit in this fiscal period

minimizing costs

Question 13 (0.5 points)

The gas station on the edge of town has gas that is 30c/gallon cheaper than my local station. I buy 10 gallons of gas a week. Assume that I value my time at $10 per hour and it takes me 1/2 hour to drive to and from the faraway station. For simplicity, assume the travel cost to the local station is zero miles and damage/depreciation of my car is trivial on the trip to the faraway station. Should I buy my gas at the faraway station?

Question 13 options:

no

yes

Rachel is assisting the CEO of her small company with a project to reduce electricity usage on the shop floor. This is a strategic engineering economic project focused on:

Question 14 options:

equipment replacement

new product introduction

cost reduction

improvement in service

Which of the following organizational structures led to massive increases in the deployment of capital in the US and UK in the nineteenth century?

Question 15 options:

Limited liability companies

Partnerships

Sole proprietorships

Worker cooperatives
Answer #1

11. opportunity cost is next best alternative foregone. 1000m ice teams or 3000 popsicles can be made.

opportunity cost for each ice cream = 3000/1000= 3

hence option 3

12. Option 1: maximising the company' value over time should be the aim of business decision makers. Short run objecticves can be anything.

13.No. I will save 0.3*10=3 dollars but I value my time for $5 for that half an hour and hence its better not to go considering opportunity cost.

14.Cost reduction option 3: As it is aimed at reducing present electricity costs. There is neither a new project or equipments or improvement in service.